# Behavior of prediction performance metrics with rare events

### v1

@@ 4 Simulation results
We display results describing the behavior of AUC estimates in Figure 1. Bias (top panel) for all algorithms decreases to zero as the effective sample size increases. The effective sample size for the AUC is driven by the number of events, i.e., for event rate , such that the three right farthest dots in each panel show the AUC bias at each event rate at a sample size of 1,000,000. We can see that within each sample size (shown with difference colors in Figure 1) as the event rate increases the bias decreases, but that once a minimum number of events is reached (approximately 1837 events), the sample size or event rate no longer impacts the bias, as bias is effectively zero. In Figure S1, we show bias and coverage versus training set size for each event rate; the trends in bias and coverage do not vary with event rate. Taken together, this shows that it is the effective sample size, not the event rate, that is driver of AUC bias. This is supported by the form of the influence function for AUC derived by LeDell et al. (2015, Theorem 4.1): the estimated AUC (and its asymptotic variance) depends on the number of events and number of non-events, and will be impacted if one of these is small.
@@
Coverage of 95% confidence intervals (bottom panel of Figure 1) based on the ESE is near the nominal level at all sample sizes (denoted with dots in Figure 1), while coverage of ASE-based intervals followed the same pattern as bias, increasing to the nominal level for all event rates and algorithms by an effective sample size of 1837. In Figure S2, we show mean squared error and confidence interval width for AUC, showing that the variance decreases with increasing effective sample size. We expect poor coverage of influence curve-based (ASE) intervals when the effective sample size is small, since these intervals use an asymptotic estimate of the variance of the cross-validated AUC and thus are asymptotic 95% intervals (LeDell et al., 2015). We observed poor behavior (bias and low coverage) for estimating AUC using logistic regression when the number of predictors exceeded (or was similar to) the effective sample size; this is consistent with work that has shown increased bias and variability in parameter estimates when the number of events per variable is low. A commonly suggested guideline, and one supported by these simulation studies, is that there should be ten or more events per variable to estimate a logistic regression model (Peduzzi et al., 1996).
@@
The performance of sensitivity and specificity are also driven by the effective sample size. Figure 2 shows sensitivity and specificity using the 95th percentile of predicted risk, while results for sensitivity and specificity at the 90th and 99th percentile following a similar pattern are found in Tables S4–S7 in the Supplementary Materials. As with AUC, the bias decreases with increasing effective sample size; coverage based on the ESE is at the nominal level for sensitivity. For specificity, while bias follows the same pattern, coverage follows a slightly different pattern. While coverage is mostly near the nominal level, when the true specificity is close to 1 (as is the case for specificity in general but particularly at the 99th percentile), there is instability of interval coverage estimates.
@@
In contrast to AUC and sensitivity, the performance of accuracy, PPV, NPV, , , and Brier score relied in part on the event rate itself, rather than just on the effective sample size. We display the results for PPV at the 95th percentile of predicted risk in Figure 3; performance for accuracy, NPV, , , and Brier score and the other algorithms is similar to the results for PPV and is provided in the Supplementary Material (Tables S8–S22, which also include all metrics at the 90th and 99th percentiles). For PPV, at a given training set size, the bias is larger for smaller event rates; this is most noticeable at the smallest training set sizes (5,000 and 10,000). Coverage was again at the nominal level. Similar results hold for the other estimands: for a given sample size, bias and coverage tended to be worse at the lower event rates (Tables S8–S22).
@@ 5 Discussion
We found through numerical experiments that the behavior (bias and confidence interval coverage) of AUC, sensitivity, and specificity is driven by the effective sample size, not the event rate. Specifically, once a minimum number of events was reached through either a larger sample size or a higher event rate, estimates of AUC and sensitivity estimates stabilized. For specificity, the effective sample size is the number of non-events and thus this stability was often reached early as we investigated rare event settings. In our simulations, as the effective sample size grew, regardless of the event rate, the cross-validated AUC converged to true AUC and empirical coverage of confidence intervals based on the influence function or the empirical standard error was close to the nominal level. Even in the setting of a very rare event—the smallest event rate we considered was , or approximately one event for every 217 non-events—these simulations show that AUC performs well (in terms of low bias for estimating the true AUC) provided that the number of events in the training set is sufficiently large. The largest training set size we considered was 1 million, which corresponds to 4,591 training set events for the smallest event rate.
While the effective sample size is a function of the event rate, it is also a function of the training set size. This implies that in large sample settings, even with a very rare event we will still have sufficiently many events for the estimated prediction metric to provide a reliable understanding of the performance of a predictive model. We do not have to worry about these prediction metrics (AUC, sensitivity, and specificity) being unreliable simply because the event rate is small. While this distinction between poor behavior of a prediction metric with respect to its reliability in rare event settings may be unimportant in settings where large data sets are not available, it is crucial in settings—such as the use of health-system electronic health records data to build clinical prediction models—where large data sets are increasingly available and computationally feasible to work with.
In contrast to AUC, the behavior of accuracy, Brier score, PPV, NPV, score, and score all showed some direct reliance on event rate in addition to being driven by the size of the training set. This is not surprising, since the marginal outcome probability is included in the definition of each of these metrics. However, bias converged to zero and coverage approached the nominal level as the effective sample size grew, regardless of event rate, indicating that the behavior of these metrics relies on both the effective sample size and the training data size.
The reliance of performance metrics on effective sample size emphasizes the need to choose an appropriate prediction modeling approach for the data. This may include dimension reduction in some cases. For example, in a situation with an effective sample size of 23 (the smallest we considered in our simulations), we would not expect a logistic regression model with 150 predictors to perform well without some dimension reduction; indeed, we observed poor performance for logistic regression in this scenario.
@@
There are two main limitations to our work. The first is that we did not vary the strength of association between outcome and predictors in our simulations: it is possible that results would be improved at a smaller effective sample size if there are strong predictors of the outcome. However, our results still provide useful information about the performance of these metrics in a scenario with complex relationships between variables. The second is that we do not have asymptotic standard errors for metrics besides AUC. While a bootstrap might be tempting in this scenario, the poor performance of empirical standard error-based intervals in some cases, due to the performance metrics being near the boundary values of 0 or 1, suggests that the standard bootstrap may not perform well in these scenarios. Bootstrap procedures designed for constructing intervals when estimates are near the boundary are available and would likely need to be implemented for these measures.
Regardless of the event rate or effective sample size, our results highlight the importance of reporting multiple prediction performance metrics and confidence intervals for these metrics, particularly when the outcome of interest is binary. In the binary outcome setting, different performance metrics (e.g., AUC, sensitivity, PPV) can provide complementary information about model behavior. Others have advocated this approach (see, e.g., Pepe et al., 2007; Hand, 2012; Lever et al., 2016; Adhikari et al., 2021). By reporting multiple metrics and describing their variability, the results can be placed in the proper context.
However, when comparing performance across prediction models and for model selection, it may be necessary to choose a single metric. This choice should be driven by careful consideration of the specific scientific interest and, when applicable, the intended use of the model. For example, if a prediction model is intended to identify patients with a 90-day risk of suicide attempt exceeding 5% (i.e., above the 95th percentile), then an important metric could be sensitivity at the 95th percentile of predicted risk. This work provides us with information on how stable and reliable we can expect our performance metric estimates to be given the event rate and sample size we have on hand and it reassures us that the effective sample size, not the event rate itself, is the factor that limits the reliability of the AUC measure.
@@
[…]

### v2

@@ 3 Results
Results describing the behavior of AUC estimates are in Figure 1. Bias (top panel) for all algorithms decreases to zero as the ESS increases. For each sample size, bias decreases as the event rate increases, but once a minimum number of events is reached (approximately 1000 events), the event rate no longer impacts bias. The three right farthest dots in each panel show AUC bias of zero at each event rate for a training sample size of 1,000,000. In Figure S1, we show that for each training set size, bias goes to zero for each event rate. These two plots show the ESS, not the event rate, drives AUC bias. This is supported by the AUC influence function form [32]: the estimated AUC (and its asymptotic variance) depends on the number of events and number of non-events, not the event rate itself.
@@
ESE-based AUC 95% CI coverage (bottom panel of Figure 1) is near the nominal level at all sample sizes, while ASE-based CI coverage followed the same pattern as bias, increasing to the nominal level by an ESS of 1000. In Figure S2, we show MSE and CI width for AUC, demonstrating the variance decreases with increasing ESS. We expect poor coverage of ASE-based intervals when the ESS is small, since these are asymptotic intervals [32]. We observed poor AUC behavior (bias and low coverage) for the logistic regression model when the number of predictors exceeded the ESS. This is consistent with work that has shown increased bias and variability in parameter estimates when the number of events per variable is low. A commonly suggested guideline, and one supported by these simulations, is there should be ten or more events per variable to estimate a logistic regression model [38].
@@
Figures 2 and 3 show that the performance of sensitivity and specificity at the 95 percentile of predicted risk are also driven by the ESS (see Supplementary Tables S4–S7 for 90 and 99 percentile, which have a similar pattern). Bias decreases with increasing ESS. ESE-based CI coverage is at the nominal level for sensitivity. For specificity, CI coverage follows a slightly different pattern; while mostly near the nominal level, when the true specificity is close to 1, there is instability in CI coverage estimates.
@@
In contrast to AUC, sensitivity, and specificity, the performance of accuracy, PPV, NPV, , , and Brier score relied in part on the event rate itself, rather than just on the ESS. We display the results for PPV at the 95 percentile of predicted risk in Figure 4; performance for accuracy, NPV, , , and Brier score and the other algorithms is similar to the results for PPV (Supplementary Tables S8–S22, which also include all metrics at the 90 and 99 percentiles). For these estimands, at a given training set size, bias and coverage tended to be worse at the lower event rates.
@@ 4 Discussion
Our numerical experiments demonstrate that bias and CI coverage of AUC, sensitivity, and specificity are driven by ESS, not event rate. Stability was reached early for specificity, as we only considered rare events and the ESS for specificity is the number of non-events. For AUC and sensitivity, once a minimum number of events (approximately 1000) was reached, through either a large sample size or higher event rate, bias was near zero and CI coverage was near nominal. Even in the smallest event rate we considered, ( approximately one event per 217 non-events) our simulations show AUC estimates had low bias as long as the training set was sufficiently large (i.e., = 400,000).
In contrast to AUC, the behavior of accuracy, Brier score, PPV, NPV, score, and score all showed some direct reliance on both event rate and training set size. This is not surprising, since the marginal outcome probability is included in the definition of these metrics. However, bias converged to zero and coverage approached the nominal level as the ESS grew, regardless of event rate. While we focused on binary classification in this article, ESS, rather than simply class proportions or number of events, likely plays an important role in bias and CI coverage of prediction performance metrics in settings with multi-class or time-to-event outcomes. Our work also emphasized the need to choose an appropriate prediction modeling approach for the data, which may include dimension reduction. For example, logistic regression did not perform well without a minimum number of events per predictor. Finally, our results highlight the importance of reporting multiple performance metrics with CIs, an approach which has been advocated by others [6, 39, 12, 25].
@@
There are some limitations to our work. First, we did not vary the strength of association between predictors and outcomes. It is possible bias could improve at smaller ESS if there are strong predictors of the outcome. Second, we do not have ASE formulas for metrics besides AUC. While the bootstrap might be tempting to construct standard errors, poor performance of ESE-based intervals, due to the performance metrics being near the boundary values of 0 or 1, suggests the standard bootstrap may not perform well. Bootstrap procedures designed for constructing CIs when estimates are near the boundary are available [40].
Our work shows that AUC estimates are stable and reliable if the ESS is large enough. That is, unbiased AUC estimates are produced if the sample size is large enough even for low event rates. Additionally, AUC ASE-based CIs had nominal coverage at large enough sample sizes, indicating that ASE-based CIs can be reliably used.
Acknowledgments
Research reported in this publication was supported by the National Institute of Mental Health of the National Institutes of Health under award numbers R01 MH125821, U19-MH099201, and U19-MH121738. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health.
@@ Acknowledgments
| • Abbreviations: GLM: logistic regression; RF: random forests; Ridge: ridge regression. | ||||||||||